OpenStudy (anonymous):

ABCD is a rectangle. If AC = 3c/9 and BD = 4-c, find the length of each diagonal.

OpenStudy (imstuck):

Can you draw the rectangle so I can see where your letters are and to what sides they belong?

OpenStudy (imstuck):

Are AC and BD diagonals?

OpenStudy (imstuck):

I bet they are, and if they are, then they are equal in length to one another. Like this: \[\frac{ 3c }{ 9 }=4-c\]

OpenStudy (anonymous):

There is no picture of a rectangle but that's what the questions looks like.

OpenStudy (imstuck):

solve for c and then fill the value you find for c back into each of the expressions.

OpenStudy (imstuck):

3c=9(4-c) 3c=36-9c 3c-9c=36 -6c=36 c=-6

OpenStudy (imstuck):

Now fill -6 in for each c in your expressions. One of them yields a -2 and the other yields a 10. Since length cannot be negative, 10 would be it.

OpenStudy (anonymous):

\\[\frac{ 3(6) }{ 9 } = \frac{ 18 }{ 9 } = 2\]

OpenStudy (anonymous):

i did it wrong oops

OpenStudy (imstuck):

That's ok. If this is a rectangle, then both diagonals are equal to one another. So use the one that gave a positive result.

OpenStudy (anonymous):

Okay thank you

OpenStudy (imstuck):

YW!!! TY for the medal!

OpenStudy (anonymous):

No problem :D